### v3

# Cleaning up the Mess: Re-Evaluating the Real-System Modeling Accuracy of Ramulator 2.0

Abstract
A MICRO 2024 best paper runner-up publication (the Mess paper [2]) with all three artifact badges awarded (including “Reproducible”) proposes a new benchmark to evaluate real and simulated memory system performance. The publication contends that Ramulator 2.0 [16] and DAMOV [5] (ZSim+Ramulator) (along with other existing memory system simulators) “poorly resemble the actual system performance” and asserts that their simulator is better.
In this paper, we show that the Mess paper [2] has 1) demonstrable technical misconfigurations, 2) methodological errors in interpreting simulation statistics, and 3) an incomplete artifact that makes its key results irreproducible. We demonstrate that the Ramulator 2.0 simulation results reported in [2] are incorrect due to multiple configuration errors instead of inherent simulation inaccuracy claimed by the Mess paper. We show that by correctly configuring Ramulator 2.0, Ramulator 2.0’s simulated memory system performance actually resembles real system characteristics well, and thus a key claimed contribution of [2] is factually incorrect. We also identify that the DAMOV simulation results in [2] use wrong simulation statistics that are unrelated to the simulated DRAM performance. By using correct DRAM simulation statistics, we show that DAMOV’s simulated DRAM latency is not constant, in contrast to [2]’s claim. Moreover, the Mess paper’s artifact repository [3, 17] lacks the necessary sources (simulator code, system configurations, memory traces, etc.) to fully reproduce all the Mess paper’s results. We find that the experiment scripts in [3, 17] use simulator executables and other resources that are neither described in the Mess paper nor found in the artifact repository [3, 17].
Our work identifies important issues in [2]’s memory simulator evaluation methodology regarding Ramulator 2.0 and DAMOV. We present results that validate the real-system modeling accuracy of Ramulator 2.0, and describe the reasons why [2]’s results with respect to these two simulators are incorrect. We emphasize the importance of carefully and rigorously validating simulation results to avoid publishing factually incorrect results and contributions. We strongly encourage the computer architecture community to consider our corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] to prevent the propagation of inaccurate and misleading results and to maintain the reliability of the scientific record. Our investigation also aims to stimulate discussion on artifact evaluation practices and on mechanisms for correcting results and artifacts after publication. To aid future works and reproduction of all our results, we open source all our code and scripts. We also discuss best practices and add sanity checks to aid users and developers of simulation tools.
11footnotetext: F. Nisa Bostancı, Haocong Luo, and Ataberk Olgun are co-primary authors.
1 Introduction
Fig. 1: Mess results for Ramulator 2.0 memory system simulation from Fig. 6 in [2] (a), Mess results for an ARM-based real system from Fig. 4 in [2] (b), our results for Ramulator 2.0 memory system simulation after correcting the errors we found in [2] (c), as explained in detail in Section 2 of this paper
Cycle-level DRAM simulators enable research focused on improving the performance, efficiency, and robustness (including security, safety, reliability) of DRAM-based memory systems by providing accurate and flexible models for DRAM and memory controller operations. Ramulator 2.0 [16, 7] (which builds on Ramulator [13, 8]), is a highly modular and extensible cycle-accurate DRAM simulator that enables rapid exploration of new ideas in DRAM-based memory systems. DAMOV [30, 5] is a simulation framework that integrates Ramulator [13, 8] into ZSim CPU simulator [32] to enable fast, scalable, and cycle-level system simulation.
A MICRO 2024 best paper runner-up publication, A Mess of Memory System Benchmarking, Simulation and Application Profiling [2], which we refer to as “the Mess paper” throughout this paper, with all three artifact badges awarded (including “Reproducible”) proposes a new benchmark to evaluate real and simulated memory system performance. While doing so, it makes some strong negative claims about Ramulator 2.0 and DAMOV, which we demonstrate to be incorrect and are due to configuration and simulator usage errors made in the Mess paper.
Fig. 1-(a) and (b) (copied from Fig. 6 and Fig. 4 in [2]) show average memory latency as a function of throughput (used memory bandwidth) for the Mess benchmark using a Ramulator 2.0 simulation and a real system, respectively, as depicted in [2]. [^1] The Mess benchmark consists of two workloads called Stream and Pointer-Chase. These workloads run concurrently. The interval between subsequent memory requests in Stream is controlled using NOPs to vary memory intensity. For example, 1K NOPs between subsequent memory requests lead to a relatively low memory intensity and 0 NOPs between subsequent memory requests lead to a very high memory intensity. Each bandwidth–latency curve in Fig. 1 is constructed by connecting Mess benchmark bandwidth–latency results in increasing memory intensity order such that the leftmost end of the curve corresponds to the smallest memory intensity and the other end of the curve corresponds to the largest memory intensity Mess benchmark configuration. Each curve depicts system performance for a unique mixture of read and write memory requests in Stream ranging from 50% reads and 50% writes to 100% reads and 0% writes. From these two figures, the authors of the Mess paper conclude that Ramulator 2.0 “poorly resembles the actual system performance” (Section I of [2]). They write “The simulated memory latency is unrealistically low and the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system” (Section IV of [2]). They attribute the discrepancy between Mess benchmark results of Ramulator 2.0 simulations and the real system to Ramulator 2.0-induced simulation errors: “This indicates that the main source of the large simulation error is indeed Ramulator 2” (Section 4 of [2]). The authors also show other unexpected results. For example, the DRAM latency results from DAMOV (ZSim+Ramulator) [5] (depicted in Section 4 of [2]) are unrealistically low and they seem to remain constant regardless of system load.
In this paper, we show that the Ramulator 2.0 simulation results in the Mess paper are incorrect and, at the time of the publication of [2], irreproducible. Fig. 1-(c) shows our reproduction of the memory bandwidth–latency curves for the Mess benchmark using the open source version of Ramulator 2.0 [16] after correcting the error we found in [2]. These results clearly show that Ramulator 2.0 does not poorly resemble real system performance contrary to what the Mess paper claims.
We carefully investigate the Mess paper’s open source artifacts [17, 3] and have communicated with the Mess paper’s authors to identify the root causes of the errors in [2]. We identify two major Ramulator 2.0 configuration errors and two other scholarship issues regarding Ramulator 2.0 in the Mess paper. We describe these issues in detail in Sections 2.4 and 2.5. In short, the Ramulator 2.0 results in the Mess paper are subject to configuration errors. Therefore, these results do not support the negative claims (asserted in the Mess paper) about Ramulator 2.0’s modeling accuracy. [^2] Strong negative assertions made in the Mess paper such as “… the most complex and trusted memory model shows the highest simulation error…” (Section 4 of [2]) are incorrect because the source of the error is the misconfiguration and misuse of Ramulator 2.0 in the Mess paper. Ramulator 2.0 can be correctly configured by careful and rigorous validation and/or by contacting Ramulator 2.0 authors and developers after observing unexpected results, to ensure that the simulator is used correctly and as intended. [^3] The Mess paper’s authors verified via email exchanges that they intentionally did not contact the developers of the memory system simulators evaluated in the Mess paper before publication of the paper.
We also identify that the reason for unrealistically low and seemingly constant DRAM latency results for DAMOV from the Mess paper is that the Mess paper’s authors report simulator statistics that are not updated during the DRAM simulation. Moreover, the Mess paper’s artifact repository [3, 17] does not contain the relevant DRAM simulation statistics from Ramulator for DAMOV. We find that the experiment scripts of the Mess paper use simulator executables and other resources that are neither described in the Mess paper nor found in the Mess paper’s artifact repository [17, 3]. We independently evaluate DAMOV’s simulated memory system performance using the open source version of DAMOV [5]. We show that, when using the correct DRAM simulation statistics from Ramulator, 1) the simulated DRAM latency does not remain constant in contrast to the results claimed by the Mess paper [2], and 2) DAMOV’s bandwidth–latency curves resemble real system characteristics.
Our work provides corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] and identifies important scholarship issues in the memory simulator evaluation methodology of the Mess paper [2]. We emphasize the importance of contacting simulator authors and developers (especially before publishing), in true open source spirit, to ensure these simulators are used with correct configurations and as intended. We strongly encourage the computer architecture community to consider our corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] to prevent the propagation of inaccurate and misleading results and to maintain the reliability of the scientific record. Our investigation also aims to stimulate discussion on artifact evaluation practices and on mechanisms for correcting results and artifacts after publication. We discuss four best practices and add sanity checks to aid users and developers of simulation tools (§6). To aid future works and reproduction of all our results, we open source all our code and scripts at https://github.com/CMU-SAFARI/Cleaning-up-the-Mess.
2 Analysis of the Evaluation of Ramulator 2.0
in the Mess Paper
2.1 The Mess Benchmark
The Mess benchmark aims to characterize a memory system based on a set of bandwidth–latency curves. To this end, the Mess paper introduces the Mess benchmark. The Mess benchmark is made up of two concurrently running workloads, Stream and Pointer Chase [35, 34]. Each curve is given by running these workloads with a pre-configured fraction of read and write requests for the Stream, ranging from 50% reads and 50% writes to 100% reads and 0% writes. Each point on a curve is given by a run of the benchmark with varying memory intensity. This is achieved by varying the interval between subsequent Stream requests by inserting NOPs. For example, 1K NOPs between subsequent memory requests lead to a relatively low memory intensity, and 0 NOPs between subsequent memory requests lead to a very high memory intensity.
2.2 Mess Paper’s [2] Ramulator 2.0 Results
Fig. 2 (Fig. 1-a plotted again for the reader’s convenience) shows the bandwidth–latency curves presented in [2] (Fig. 6 in [2]) for Ramulator 2.0 obtained with the Mess benchmark. This result unexpectedly and oddly shows that the simulated memory system 1) uses significantly smaller bandwidth than what is available (depicted as the “Max. theoretical BW” in the figure) and 2) yields a relatively small memory access latency.
Fig. 2: Bandwidth–latency curves for Ramulator 2.0 (copied from Fig. 6 in [2])
2.3 Reproducibility Issues in [2]
We note that it was not possible to reproduce any Ramulator 2.0 result from the Mess paper artifacts available at the time of the publication and artifact evaluation of the Mess paper [17, 3]. Thus, we contacted the authors of [2] on 16 October 2024 to obtain 1) the Ramulator 2.0 source code and 2) the Mess benchmark memory traces used in Ramulator 2.0 simulations reported in [2]. The missing sources needed to replicate the Mess paper’s Ramulator 2.0 results were eventually uploaded to the Mess GitHub repository [17] following our contact on 18 July 2025, 10 months after the Mess paper was awarded (on 8 September 2024) all three artifact evaluation badges.
2.4 Fundamental Configuration Errors in [2]
We identify two fundamental configuration errors in the Mess paper’s Ramulator 2.0 simulations. We attribute the discrepancy in bandwidth and memory access latency between the Mess paper’s Ramulator 2.0 simulations and real system evaluation to these fundamental errors. We also identify two other scholarship issues.
First, the Mess paper’s scripts (in the newly-uploaded source code on 18 July 2025) instantiate and use a single-channel DDR5 Ramulator 2.0 memory system. The authors extrapolate single-channel DDR5 bandwidth measurements to 16-channel DDR5 bandwidth measurements by multiplying the observed bandwidth by 8 instead of 16. This is because they incorrectly assume that a DDR5 channel in Ramulator 2.0 is 64 bits wide, consisting of two independent 32-bit channels. However, the source code of Ramulator 2.0 clearly shows that a DDR5 channel is 32 bits wide [^4] The m_channel_width parameter in DDR5.cpp in the Ramulator 2.0 repository [7]. in Ramulator 2.0, which is consistent with JEDEC DDR5 standards [10, 12, 11]. [^5] The JEDEC DDR5 SDRAM component [10] and UDIMM module [12] standards consistently use the term “channel”. The JEDEC DDR5 RDIMM standard [11] uses the terms “channel” and “sub-channel” interchangeably. However, both “channel” and “sub-channel” in these JEDEC DDR5 documents [10, 12, 11] are 32-bit wide. The Mess paper misleadingly presents (extrapolated) 8-channel (32-bit each) DDR5 results as 16-channel (32-bit each) DDR5 results, leading to incorrect claims like “…the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system” in Section 4 of [2]. The maximum bandwidth of an 8-channel DDR5-4800 [10] is 153.6 GB/s [10], not 307 GB/s. Thus, a bandwidth of more than 153.6 GB/s should not be expected of an 8-channel DDR5-4800 configuration. We show that a correctly configured 16-channel DDR5-4800 Ramulator 2.0 simulation achieves 281.1 GB/s bandwidth (as depicted in Fig. 1-c; see Section 4 for a detailed description).
Second, the Mess paper’s scripts use the SimpleO3 frontend of Ramulator 2.0 [7] with unrealistically low latency configurations compared to a real system. For example, they configure 1) the cache latency to be 0 CPU clock cycles and 2) each CPU core to have 1024 miss-status holding registers (MSHRs). Compared to a real CPU, such a configuration has 1) significantly smaller round-trip latency between a core and DRAM, and 2) significantly lower queuing delay for memory requests when memory intensity is high. About Ramulator 2.0’s simulation accuracy, the Mess paper claims that “…the simulated memory latency is unrealistically low…” (Section 4 of [2]). Because the unrealistic cache setup is in part responsible for the low latency observed in [2], the cache setup in [2] constitutes a fundamental configuration error. Thus, the Ramulator 2.0 simulation accuracy claim made in the Mess paper [2] is incorrect.
2.5 Scholarship Issues in [2]
First, the Mess paper does not disclose the unrealistically low latency configurations of the SimpleO3 frontend of Ramulator 2.0 in the paper. The Ramulator 2.0 configuration file available in the artifact repository also does not include any of these configurations. After the Mess paper’s authors uploaded their Ramulator 2.0 source code, we found (via email exchanges with the Mess paper’s authors) that they implement the unrealistically low latency configurations by directly modifying the source code of Ramulator 2.0, without any documented changes.
Second, the Mess paper uses different Mess benchmark methodologies for real systems and simulators without disclosing this difference in the paper. [^6] Identified via email exchanges with the Mess paper’s authors. For real systems (e.g., Fig. 5-a in [2]), the Mess benchmark is run as described in the Mess paper (i.e., by concurrently running pointer-chasing and Stream workloads and measuring the memory latency of the pointer-chasing workload’s memory accesses). For simulators (e.g., Figures 5-b, -c, -d, -e, and -f in [2]), the authors do not run a pointer-chasing workload and they measure the average memory latency for only the Stream workload. Thus, the Mess paper’s authors compare apples to oranges by running and measuring different workloads in simulators and in real systems. And, this difference in simulation and real system evaluation methodologies is not disclosed in the Mess paper.
Through our comprehensive investigation of [2]’s Ramulator 2.0 artifacts, we conclude that the Ramulator 2.0 results in [2] are subject to multiple configuration and methodological errors that could have been avoided by careful and rigorous validation.
3 Our Evaluation Methodology
To show that Ramulator 2.0, when configured properly and used correctly, yields reasonable and realistic bandwidth–latency curves, we evaluate a modern DDR5 memory system in Ramulator 2.0 using the same access patterns as described in the Mess benchmark.
We introduce a new Ramulator 2.0 frontend module, the Mess Request Generator, which we open source, along with all our code and scripts in our repository [6, 4] dedicated to this work. Our Mess Request Generator frontend is designed to emulate the memory access patterns of the Mess benchmark by generating a mixture of random access (random) and streaming (stream) read requests. The request generator issues random requests sequentially (to emulate a pointer-chasing workload), meaning that the request generator issues a new random request only when the previous random request returns from the memory controller. The request generator issues a configurable mix of stream read and stream write requests to the memory controller as often as every DRAM command interface clock cycle (e.g., $0.416\text{\,}\mathrm{ns}$ for DDR5-4800 [10]). We modulate the frequency of stream requests using NOPs between two consecutive requests. The sequence of stream requests are designed to take advantage of bank-group- and bank-level parallelism and row hits to fully utilize DDR5 data transfer bandwidth. To do so, we use a stream access pattern that overlaps the latency of memory requests by issuing requests that are interleaved across bank-groups and banks in each channel. The request generator keeps issuing stream requests while waiting for a random read request to return from the memory controller.
Read/Write Mix. The Mess Request Generator frontend for Ramulator 2.0 is configurable with a read ratio between 0.5 and 1, which is applied to the stream memory requests. A read ratio of 1 corresponds to streaming accesses consisting of 100% reads, while a read ratio of 0.5 yields 50% read and 50% write requests for the streaming memory accesses. We use read ratios of 0.5, 0.6, 0.7, 0.8, 0.9, and 1.0 in our evaluation.
Stream Bandwidth Use Modulation. To simulate various system loads, we further extend the Mess Request Generator with a configurable NOP frequency. In every Ramulator 2.0 frontend cycle, the Mess Request Generator can either issue a memory request or not, depending on the value of the NOP frequency. For example, with a NOP frequency of 1, every cycle leads to a memory request being issued, while a NOP frequency of 100 means that every 100 simulated cycles, a memory request is issued.
We use the Mess Request Generator frontend for Ramulator 2.0 to construct the bandwidth–latency curves. We use the Ramulator 2.0 memory system configuration in Table 1. We open source our Ramulator 2.0 configuration along with all source code and scripts to reproduce our results in [6, 4]. [^7] Our Ramulator 2.0 codebase [6, 4] dedicated for this work introduces only minimal and necessary changes to the open source Ramulator 2.0 repository [7]. First, we implement the Mess Request Generator frontend we describe. Second, we add the timing parameters for the DDR5-4800AN standard [10] to remain consistent with what is evaluated in the Mess paper [2]. Third, we disable write-to-read forwarding in the memory controller to ensure all memory requests are served by DRAM. Fourth, we add utility functions to record various statistics about memory requests and to programmatically configure the Mess Request Generator.
Table 1: Simulated System Configuration
Frontend: Mess Request Generator for Ramulator 2.0 [16]
DRAM: DDR5-4800AN [10]; 16 channels (32-bit each), 1 rank, 8 bank groups, 4 banks
[…]
Memory System: FR-FCFS request scheduling policy [31, 36], All-bank refresh, Read queue size: 32, Write queue size: 32
To construct each bandwidth–latency curve, we run the simulation for 20K random read requests and measure only the latency of the random read requests (i.e., same as in the Mess benchmark). We sweep NOP frequency from 1 to 10000 to vary system load.
4 Key Results
[…]
Taking such refresh operations and their impact on data transfers into account, we compute the maximum achievable bandwidth from the maximum theoretical bandwidth as 281.2 GB/s (indicated by the green line in the figure) from the following equations:
$BW_{achievable}=BW_{theoretical}\cdot\left(1-\frac{REF_{penalty}}{tREFI}\right)$
$REF_{penalty}=tRTP+tRP+tRFC+tRCD$
Fig. 3: Bandwidth–latency curves for DDR5-4800AN in Ramulator 2.0 using the Mess Request Generator frontend and configuration in Table 1
We make three major observations from Fig. 3. First, we observe that the maximum achieved memory bandwidth (281.1 GB/s for a read ratio of 1, i.e., 100% reads, and NOP frequency of 1, i.e., no NOP cycles) almost exactly matches the maximum achievable memory bandwidth. We thoroughly investigate the DRAM command sequence modeled by Ramulator 2.0 to understand the small difference between the maximum achieved bandwidth (281.1 GB/s) and the maximum achievable bandwidth (281.2 GB/s) values. We observe that the $REF_{penalty}$ slightly varies for each periodic refresh operation. For example, if the memory controller opens a DRAM row to serve a random access read request and schedules a periodic refresh operation immediately after this read request, the delay between two $READ$ commands interrupted by a refresh operation is slightly higher than $tRTP+tRP+tRC+tRCD$: (tRAStRCD)+tRP+tRC+tRCD(tRAS-tRCD)+tRP+tRC+tRCD.
Second, as the fraction of write memory requests increases, the bandwidth–latency curves are curved more aggressively towards the top left corner of the figure because a larger fraction of write memory requests induces a higher latency and bandwidth penalty due to more frequently incurred DRAM bus turnaround latency [15].
Third, we observe that the latency of the random access memory requests consistently increases with used memory bandwidth because interference from stream requests in the memory controller read/write queues increases, which in turn increases the queuing delay for the random read requests. [^8] Note that the memory requests are only subject to delays incurred at the memory controller because our simulation injects read and write requests directly into the memory controller read and write queues.
We conclude that Ramulator 2.0, when configured and used correctly, yields reasonable and realistic bandwidth–latency curves that resemble the ones obtained from real systems by [2] (see Section 3, Fig. 3 in [2]). The results we present in this section thus show that the Ramulator 2.0 results presented in [2] are incorrect and are not representative of Ramulator 2.0’s modeling accuracy and capability.
5 Evaluation of DAMOV in the Mess Paper
5.1 Mess Paper’s [2] DAMOV (ZSim+Ramulator) Results
Fig. 4 shows the bandwidth–latency curves of various ZSim [32] memory models (sub-figures b, c, d, e, and f), compared to real system measurements (sub-figure a), from the Mess paper [2] (Fig. 5 in [2]). The Mess paper claims DAMOV (i.e., ZSim+Ramulator) [5] “provides a fixed 25 ns latency in the whole bandwidth area and for all memory traffic configurations” (Section 4 in [2]).
Fig. 4: Bandwidth–latency curves for ZSim memory models from the Mess paper [2] (same as Fig. 5 in [2])
To understand why the Mess paper shows such an unrealistic (i.e., low and seemingly constant) memory latency for DAMOV, we carefully checked the code and scripts from the artifact repository of the Mess paper [3, 17, 18]. We identify two major issues: 1) irreproducible simulation, and 2) erroneous memory latency statistics usage, which we describe in Sections 5.2 and 5.3.
5.2 Reproducibility Issues in [2]
We carefully checked the artifact repositories of the Mess paper[17, 3, 18] based on the description of the artifact in Appendix C of [2]. Unfortunately, we could not find the following three items used by the Mess paper to produce the results in Fig. 5 (f) in the Mess paper: 1) the source code of the ZSim simulator, 2) the source code of Ramulator, and 3) the configuration of Ramulator used. We also checked the path to the simulator executable in the experiment scripts from the Mess paper artifact [3, 17, 18]. Unfortunately, the hardcoded path999/home/bsc18/bsc18278/zsimdramsim3/acmSimulation/zsim.git_mt/build/debug/zsim from line 33 of Mess-benchmark/CPU/Simulators/Execution-driven/ZSim/ramulator/submit.batch in [17, 3] points to a directory that does not exist in the artifacts [17, 3, 18].
We tried to independently reproduce the Mess paper’s results (i.e., Fig. 5-f in [2]) by combining the ZSim configuration available in the artifact repository of the Mess paper with the Ramulator configuration from DAMOV [5]. We modify the Ramulator configuration to model the $6\timesDDR$4-2666 configuration from the Mess paper in a best-guess manner. We perform the simulation using 1) the ZSim and Ramulator code from DAMOV [5], and 2) the Mess paper’s stream_mpi workload. However, doing so triggers a rigid assertion in DAMOV because it does not support non-power-of-2 numbers of DRAM channels. Therefore, we cannot independently reproduce the DAMOV results in Fig. 5-f of the Mess paper using the configuration that the Mess paper describes that it used to produce Fig. 5-f in [2].
5.3 Technical Simulator Usage Errors in [2]
We find that the Mess paper uses two statistics from ZSim to calculate the average memory latency for DAMOV, latGETnl (i.e., the accumulative latency) and mGETs (i.e., the number of GET requests) from the L1-D cache. [^10] From lines 162-163 in Mess-benchmark/CPU/Simulators/Execution-driven/ZSim/ramulator/processing/calculator.py These are not the correct statistics to compute the memory latency for DAMOV because the bound-weave simulation methodology of ZSim [32] 1) only updates latGETnl with constant latencies (i.e., contention-free latency) throughout the memory hierarchy during the bound phase, and 2) only updates the core clock cycle (instead of latGETnl) with the actual simulated DRAM latency from Ramulator during the weave phase. In other words, the latGETnl statistics from ZSim does not include any actual DRAM latency simulated by Ramulator, but only includes the constant contention-free latency.
We find that the average memory latency for DAMOV as the Mess paper measures (i.e., latGETnl / mGETs) is about 48 cycles. This matches closely with the accumulative contention-free latency from L1-D to the memory controller of 52 cycles (i.e., about 25ns for 2.1GHz CPU frequency) based on the ZSim configuration [^11] From Mess-benchmark/CPU/Simulators/Execution-driven/ZSim/ramulator/sb.cfg of the Mess paper [2].
The DRAM performance metrics should come from Ramulator statistics output itself. Unfortunately, all the Ramulator statistics output [^12] E.g., Mess-benchmark/CPU/Simulators/Execution-driven/ZSim/ramulator/measuring/bw-lat/measurment_100_0/bwLatCurves.ramulator.stats in the Mess artifact are empty. We therefore believe that the results in the Mess paper for DAMOV are reflecting incorrect statistics due to simulator usage error. Our findings demonstrate the need to 1) carefully configure, understand, and use simulators, and 2) validate findings before making strong claims about the accuracy of the simulators.
5.4 Our Evaluation Methodology
To show that DAMOV (ZSim+Ramulator), when used correctly, yields reasonable bandwidth–latency curves and non-constant DRAM latency, we evaluate a modern DDR4 memory system in DAMOV using the Mess benchmark’s streaming access workload and the ZSim configuration files in the Mess paper’s artifact [17]. Table 2 describes our system simulation configuration. We make three modifications to the configuration. First, to bypass the assertion failure that is described in Section 5.2, we instantiate DRAM with 8 channels. This configuration matches the memory system of another real DDR4-based system that the Mess paper evaluates, IBM Power9 [2]. Second, we add the timing parameters of the DDR4-2666T standard [9] to remain consistent with what is evaluated in the Mess paper [2]. Third, we revise the Mess paper’s configuration file to avoid runtime errors (caused by configurations that DAMOV’s open source version does not support [5]).
To construct each bandwidth–latency curve, we run the simulation for 100 million instructions. We sweep NOP frequency from 1 to 750 and the read ratio from 50 to 100 to vary system load. We report the memory access latency for read requests using Ramulator statistics. This enables us to measure the memory access latency at the memory controller level, excluding the delays caused by the core model and the cache hierarchy. [^13] To measure the end-to-end latency of a memory access, the ZSim core model needs to be modified to collect this statistic.
Table 2: Simulated System Configuration in DAMOV
Processor: 24 cores @ 2.4 GHz
Cache Hierarchy: L1i and L1d: 8-way 64 KB, L2: 16-way 1 MB, LLC: 11-way 33 MBs.
DRAM: DDR4-2666T [9]; 8 channels (64-bit each), 1 rank, 4 bank groups, 4 banks
[…]
Memory System: FR-FCFS request scheduling policy [31, 36], All-bank refresh, Read queue size: 32, Write queue size: 32
5.5 Key Results
Fig. 5 shows the bandwidth–latency curves generated using DAMOV with the configuration given by Table 2. The maximum theoretical bandwidth for 8 channels DDR4-2666T (indicated by the orange line in the figure) is 170.6 GB/s. The maximum bandwidth cannot be sustained during standard operation due to all-bank periodic refresh operations [9] that conflict with and delay data transfers. Taking such refresh operations and their impact on data transfers into account, we compute the maximum achievable bandwidth from the maximum theoretical bandwidth as 164.2 GB/s (indicated by the green line in the figure) from the $BW_{achievable}$ equation in §4.
Fig. 5: Bandwidth–latency curves for DDR4-2666T in DAMOV using the configuration in Table 2.
From Fig. 5, we observe that DAMOV’s memory access latency is not constant, in contrast to what is reported in the Mess paper. As system load increases, the memory access latency increases. We conclude that when correct DRAM statistics are used, DAMOV’s simulated memory access latency is not constant.
6 Discussion
For Ramulator 2.0 and DAMOV users to avoid the configuration and usage errors described in detail in this work, we provide four best practices for Ramulator 2.0 and DAMOV, and we integrate memory bandwidth statistics as a sanity check into our open sourced version of Ramulator 2.0.
6.1 Ramulator 2.0 and DAMOV Best Practices
Best Practice 1. We recommend that Ramulator 2.0 users familiarize themselves with JEDEC standard documents (e.g., the JEDEC DDR5 DRAM standard [10]) instead of relying on information other than the official standard or the simulator documentation.
Importantly, users should be aware that colloquial technical terms may have different meanings compared to precise definitions in official standards and simulator implementations.
Best Practice 2. Users should understand Ramulator 2.0’s memory organization and configurations, which are documented in the code and the repository.
The first Ramulator 2.0 configuration error (see 2.4) in the Mess paper [2] could have been avoided with the first two best practices.
Best Practice 3. Users should understand simulation methodology and simulator implementations sufficiently such that they use the correct statistics.
The third best practice could have prevented DAMOV simulator usage errors (see 5.3) in [2].
Best Practice 4. To ensure correct simulator usage, we recommend communicating with the simulator developers or user community when unexpected results are observed (in line with the open source spirit of free software development [14]).
All configuration and usage issues in [2] that are described in our paper could have been avoided with the fourth best practice.
6.2 Memory Bandwidth Statistics as a Sanity Check
Ramulator 2.0 outputs standard statistics that are useful for performance evaluation and debugging. To allow Ramulator 2.0 users to more easily determine if their simulation is acting unexpectedly with respect to the maximum and used memory bandwidth, we add two new statistics to our open sourced version of Ramulator 2.0. First, the maximum bandwidth statistic quantifies the maximum theoretical memory bandwidth that can be attained by the simulated DRAM configuration ($BW_{theoretical}$ in §4). Second, the used memory bandwidth statistic quantifies the DRAM throughput attained by the simulated workload (e.g., a point along a curve in Fig. 3).
7 Conclusion
We demonstrated that the claims made by the Mess paper about Ramulator 2.0 [16] and DAMOV [5] are incorrect and are due to simulator configuration and simulator usage errors made in the Mess paper [2]. We strongly encourage the computer architecture community to consider our corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] to prevent the propagation of inaccurate and misleading results and to maintain the reliability of the scientific record. More broadly, simulators should be configured, understood, and used very carefully and rigorously, especially before making claims about them being wrong. The open source spirit also suggests that the authors of freely available simulators are properly contacted to ensure these simulators are used with correct configurations and as intended, especially before publishing strong and incorrect claims about such simulators. Our investigation also aims to stimulate discussion on artifact evaluation practices and on mechanisms for correcting results and artifacts after publication. To facilitate transparency and reproducibility, we open source all our code and results at https://github.com/CMU-SAFARI/Cleaning-up-the-Mess.
Acknowledgments
We thank the anonymous reviewers of ISPASS 2026 for feedback. The anonymous reviewers heavily scrutinized this paper and provided numerous suggestions, in some cases sentence by sentence, to ensure that each statement in this paper is correct and firmly backed up by data. If the strict scrutiny applied to this paper had been consistent across the scientific review process, we believe the Mess paper would not have been published in its current state. We believe our results and analyses in this paper open up broader questions regarding the review and artifact evaluation processes. In particular, it is critical for tool papers that claim incorrect results in freely available open-source software to be held to the standard of being asked to have checked with the authors and maintainers of the software first, to establish the accuracy of their results. We thank practitioners from many major companies in industry and researchers at many institutions in academia who have been using Ramulator and providing valuable feedback since the release of the first version in 2015, as well as the second version in 2023. Their experiences and comments helped us validate the practicality, usability, and accuracy of our tools. Some of these folks have also read the Mess paper and expressed their surprise that it was selected as a "best paper runner up" and even accepted, even though its authors had no communication with the developers and maintainers of Ramulator 2.0. We thank the SAFARI Research Group members for constructive feedback and the stimulating & open intellectual and scientific environment they provide. We acknowledge the generous gift funding provided by our industrial partners (especially Google, Huawei, Intel, Microsoft), which has been instrumental in enabling the extensive research we have been conducting on memory systems for more than two decades [19, 26, 24, 21, 23, 20, 1, 25, 33, 27, 28, 29, 22]. This work was in part supported by a Google Security and Privacy Research Award and the Microsoft Swiss Joint Research Center.
Appendix A Artifact Appendix
A.1 Abstract
This artifact provides the source code and scripts to reproduce the key experiments and their results presented in our ISPASS 2026 paper. This artifact enables reproducing the following key results and figures:
  1. 1.
    Bandwidth–latency curves for DDR5-4800AN in Ramulator 2.0 (Figures 1.c and 3)
  2. 2.
    Bandwidth–latency curves for DDR4-2666T in DAMOV (Figure 5)
The artifact is built on two simulation environments: Ramulator 2.0 and DAMOV. These two simulation environments and the corresponding experiments are independent, and we intend them to run separately.
A.2 Artifact Check-List (Meta-Information)
  • Program: C++ programs, Python scripts, shell scripts.
  • Compilation: clang, cmake, GNU make, scons.
  • Run-time environment: Linux (tested on Ubuntu 20.04 and 22.04), Python 3, (optional) Docker.
  • Execution: Python scripts.
  • Metrics: Memory access latency (in ns), Used memory bandwidth (GB/s).
  • Output: Figures in PDF and PNG format and related results as script outputs.
  • How much disk space required (approximately)?: 5 GB.
  • How much time is needed to prepare workflow (approximately)?: 10-30 minutes.
  • How much time is needed to complete experiments (approximately)?: 1-2 hours, given enough compute resources (e.g., approximately 200 CPU cores).
  • Publicly available?: Yes.
  • Archived (provide DOI)?: Yes, DOI: https://doi.org/10.5281/zenodo.18911020
A.3 Description
A.3.1 How to Access
The source code and scripts can be downloaded from Zenodo (https://doi.org/10.5281/zenodo.18911020).
A.3.2 Hardware Dependencies
The artifact is designed for Linux-based systems and can be run either on a Slurm-based cluster infrastructure or on a personal computer.
The full evaluation of our DAMOV results requires many multithreaded simulation runs, therefore, we strongly recommend using a Slurm-based environment with many CPU cores whenever possible. To enable easy reproduction of our DAMOV results, we provide SSH access to our internal Slurm-based infrastructure, including all required hardware and software, during the artifact evaluation. Please contact us through HotCRP and/or the AE committee for details.
Our tested system specifications:
  • Docker environments with Linux 22.04 images assuming x86-64 systems
  • (For DAMOV experiments) Supporting at least 48 GB memory and 24 threads per experiment
  • (For Ramulator 2.0 experiments) Supporting at least 6 GB memory per single-threaded experiment
A.3.3 Software Dependencies
  • Linux Operating System (tested on Ubuntu 20.04 and 22.04).
  • (Optional) Docker
  • (Recommended) Slurm
  • Python3 (version 3.8)
  • Packages: pandas numpy matplotlib seaborn pyyaml scipy scons
  • CMake, GNU Make
All experiments are run with Python scripts:
apt-get install python3-dev python3-pip
Plotting packages can be installed with:
pip3 install matplotlib pandas numpy
All requirements are installed with build scripts automatically:
  • DAMOV: build-essential, scons, automake, autoconf, m4, perl, flex, bison, byacc, libhdf5-dev, libelf-dev, python3-dev, python3-pip, openmpi-bin, libopenmpi-dev.
  • Ramulator 2.0: argparse, spdlog, yaml-cpp.
A.4 Installation and Experiment Workflow
A.4.1 Ramulator 2.0 Experiments
Ramulator 2.0 experiments are intended to run on a personal computer using a Docker image that handles all dependencies and sets up the simulation environment.
To build the simulator, run all experiments and plot inside the Docker container:
cd ramulator2
./build_in_docker.sh
This will build ramulator2 from source inside an Ubuntu 22.04 container, run all simulation experiments (run.py), generate the plots (plot.py), and copy the figures (latency_bandwidth_16ch.pdf and latency_bandwidth_16ch.png) and figures back to the mess/ directory.
Alternatively, you can build and run Ramulator 2.0 natively by following the instructions in the ramulator2/README.md file.
Customizing Evaluation Setup. You can configure the maximum number of threads to concurrently run the experiments in ramulator2/mess/run.py
# Number of concurrent threads
MAX_WORKERS = 16
A.4.2 DAMOV Experiments
DAMOV experiments are intended to run on a Slurm-based infrastructure using a native build or a Docker image. Our scripts satisfy all requirements and run experiments automatically.
Note that during the dependency setup phase, the script first tries to set up libraries with apt-get command, which can yield error outputs when the user does not have the privilege to execute this command. If this happens, the script falls back to setting up libraries locally by downloading the packages and copying them to the appropriate directories (/simulator/libconfig).
To set up dependencies, compile simulators and workloads, and run experiments in native environment:
cd DAMOV/
python3 run_artifact.py
(Optional) To use a Docker container:
cd DAMOV/
python3 run_artifact_container.py
This will build DAMOV and workloads from source, and schedule 700+ Slurm jobs.
(Optional) For a fast verification with fewer experiments:
# native:
python3 run_artifact.py –fast
# Docker container
python3 run_artifact_container.py –fast
This will build DAMOV and workloads from source, and schedule fewer Slurm jobs (e.g., 10-20).
Customizing Evaluation Setup. Please see the DAMOV/README.md file for instructions to further customize the evaluation setup.
Plotting the Results After all experiments are complete, you can generate Figure 5 from the paper.
python3
simulator/scripts/plot_latency_bw_curve.py
This will generate the figures (figure5.png and figure5.pdf) in the DAMOV/ directory.
A.5 Evaluation and Expected Results
A.5.1 Ramulator 2.0 Results
Figure 1.c and 3 are expected to be reproduced as ramulator2/mess/latency_bandwidth_16ch.pdf. We expect the run script to reproduce the two figures and output the maximum bandwidth reported in our paper.
A.5.2 DAMOV Results
Figure 5 is expected to be reproduced as DAMOV/figure5.pdf. We expect the run script to reproduce the figure and output the maximum bandwidth reported in our paper. Please note that DAMOV experiments include multithreaded experiments, which result in small changes in the produced curves and the maximum bandwidth value across experiment runs. However, the conclusions of our paper still apply to the reproduced figures and results, as the trends remain the same.
A.6 Troubleshooting
To retry failing Slurm jobs (for any reason), the run script (e.g., DAMOV/run_artifact.py) can be run again. By default, the script skips completed experiments and only resubmit/rerun an experiment if it failed previously. Note that the run script should be executed after all remaining jobs are completed to avoid disrupting existing jobs.
A.7 Methodology
Submission, reviewing, and badging methodology:
Appendix B Evaluating the Mess Benchmark
on Standalone Ramulator
The Mess paper [2] evaluates the memory access latency and the used memory bandwidth behavior of DAMOV (i.e., ZSim+Ramulator) running the Mess benchmark (Fig. 5 in [2]) and claims that DAMOV "provides a fixed 25 ns latency in the whole bandwidth area and for all memory traffic configurations" (Section 4 in [2]). After careful analysis of the artifact repositories of the Mess paper [2, 17, 3], we find that the Mess paper uses two L1-D statistics (i.e., latGETnl and mGETs) from ZSim to calculate the average memory latency for DAMOV, and not the correct DRAM statistics provided by Ramulator. We independently evaluate DAMOV’s simulation accuracy in §5.5 and show that when correct simulation statistics are used, DAMOV’s memory access latency is not constant (Fig. 5 in §5.5).
Our results from this investigation (Fig. 5 in §5.5) depict a gap between the maximum used memory bandwidth (shown as the rightmost endpoint of the red curves) and the maximum achievable bandwidth (shown as the green line). In this section, we investigate this bandwidth gap to better understand the underlying simulated behavior better.
We hypothesize that the bandwidth gap is caused by a combination of workload characteristics, system configuration, and the integration of the core and memory models. DAMOV executes the Mess benchmark by compiling and running the benchmark resources provided in the Mess paper’s artifact [17]. This is inherently different from crafting the memory access pattern of the Mess benchmark with a dedicated frontend module (similar to our Ramulator 2.0 evaluation methodology described in §3). The memory access pattern of the workload executed on the simulated core model depends on many factors that might specifically hurt the streaming access pattern (e.g., DRAM address mapping) or yield a memory access frequency that is not sufficiently stressing the memory (e.g., core model optimizations). Therefore, this setup may not yield a memory access pattern that fully utilizes all available memory bandwidth.
To determine the cause of this gap and whether it is due to the memory model of Ramulator [13, 8], we evaluate a DDR4 memory system in Ramulator using the same access pattern as described in the Mess benchmark. Note that the Ramulator version used in DAMOV is not Ramulator 2.0 and is an earlier Ramulator version open-sourced at [8]. Similar to our evaluation methodology for our Ramulator 2.0 results (§3), we introduce a new Ramulator frontend module, called MessProcessor. The MessProcessor frontend is designed to emulate the memory access patterns of the Mess benchmark by generating a mixture of random access and streaming read requests, similar to the Mess Request Generator for Ramulator 2.0. Using the MessProcessor frontend, standalone Ramulator performs memory accesses with the patterns of the Mess benchmark generated directly within Ramulator. In contrast, DAMOV executes the Mess benchmark (from the artifact repository of the Mess paper [4]), and the memory requests are generated by the ZSim cores and fed to Ramulator’s memory system.
We use the MessProcessor frontend to construct the bandwidth–latency curves for a DDR4 system. We use the Ramulator memory system configuration in Table 3. We open source our Ramulator configuration, along with all source code and scripts to reproduce our results [4].
Table 3: Simulated System Configuration for Standalone Ramulator
Frontend: MessProcessor for Ramulator [13].
DRAM: DDR4-2666T [9]; 8 channels (64-bit each), 1 rank, 4 bank groups, 4 banks
[…]
Memory System: FR-FCFS request scheduling policy [31, 36], All-bank refresh, Read queue size: 32, Write queue size: 32
Fig. 6 shows the bandwidth–latency curves generated using standalone Ramulator with the MessProcessor frontend and the configuration given by Table 3. The maximum theoretical bandwidth (yellow line) and the maximum achievable bandwidth (green line) are calculated as described in §5.5.
Fig. 6: Bandwidth–latency curves for DDR4-2666T in standalone Ramulator [13, 8] using the MessProcessor frontend and configuration in Table 3
We make two key observations from Fig. 6. First, the maximum used memory bandwidth (158.9 GB/s for a read ratio of 1, i.e., 100% reads, and NOP frequency of 1, i.e., no NOP cycles) is close to the maximum achievable bandwidth (164.2 GB/s). We attribute the remaining difference between the values to the random accesses in the Mess benchmark access pattern. Achieving maximum bandwidth requires exploiting row buffer locality and all levels of parallelism (e.g., bank-level and bank-group-level). Random accesses in the Mess benchmark access pattern can introduce row buffer conflicts and reduce the bank and bank-group-level parallelism.
Second, the maximum used bandwidth reported by the standalone Ramulator (with the MessProcessor frontend) is higher than the maximum used bandwidth reported by DAMOV (i.e., ZSim+Ramulator). This indicates that the bandwidth gap between DAMOV’s maximum used bandwidth and the maximum achievable bandwidth is not caused by the memory model of Ramulator.
We conclude that the standalone Ramulator yields reasonable bandwidth–latency curves with a higher maximum used bandwidth value compared to DAMOV, and thus, the bandwidth gap between the maximum used bandwidth and the maximum achievable bandwidth is not due to Ramulator’s memory model.